# When the Rules Change: Adaptive Signal Extraction via Kalman Filtering and Markov-Switching Regimes

### v1

@@ Abstract
Static linear models of order flow assume constant parameters, failing precisely when they are needed most: during periods of market stress and structural change. This paper proposes a dynamic, state-dependent framework for order flow signal extraction that adapts to shifting market conditions in the Korean stock market. Using daily transaction data from 2020–2024 covering 2,439 stocks and 2.79 million stock-day observations, we implement three complementary methodologies: (1) an Adaptive Kalman Filter where measurement noise variance is explicitly coupled to market volatility; (2) a three-state Markov-Switching model identifying Bull, Normal, and Crisis regimes; and (3) an Asymmetric Response Function capturing differential investor reactions to positive versus negative shocks. We find that foreign investor predictive power increases 8.9-fold during crisis periods relative to bull markets ($\beta_{crisis}=0.00204$ vs. $\beta_{bull}=0.00023$), while individual investors exhibit momentum-chasing behavior with 6.3 times stronger response to positive shocks. The integrated “All-Weather” strategy provides modest drawdown reduction during extreme market events, though challenges remain in the post-COVID high-rate environment.
@@ 1 Introduction
This paper addresses this limitation by developing a dynamic, state-dependent framework for order flow analysis that explicitly adapts to changing market conditions. Our approach rests on three methodological pillars:
First, we implement an Adaptive Kalman Filter that treats the “true informed signal” as a hidden state obscured by noise trading. Unlike standard implementations, we make measurement noise variance heteroskedastic, explicitly coupling it to realized market volatility. When volatility spikes, the Kalman gain automatically decreases, causing the model to discount high-volume panic trading as noise rather than signal.
Second, we employ a Three-State Markov-Switching Model to identify discrete market regimes—Bull, Normal, and Crisis—each with distinct return dynamics and order flow relationships. This allows parameters to “snap” to new equilibria rather than merely drift, capturing the qualitative changes in market behavior during stress periods.
Third, we estimate Asymmetric Response Functions that capture how different investor types react to positive versus negative shocks. This reveals behavioral heterogeneity that static models obscure: foreign investors act as contrarian liquidity providers during crashes, while retail investors engage in momentum-chasing behavior that amplifies market movements.
Using daily transaction data from the Korean stock market spanning 2020–2024, we document several key findings. Foreign investor order flow exhibits dramatically higher predictive power during crisis regimes ($\beta=0.00204$, $t=19.23$) compared to bull markets ($\beta=0.00023$, t=7.81t=7.81)—an 8.9-fold increase. Individual investors display asymmetric responses to market shocks, with response coefficients to positive shocks ($\beta^{+}=0.000089$) far exceeding those to negative shocks ($\beta^{-}=0.000014$), consistent with momentum-chasing behavior. The integrated strategy incorporating all three methodologies provides modest improvement in maximum drawdown during crisis periods, though overall sample performance remains challenging in the post-COVID environment.
Our contribution is methodological and empirical. Methodologically, we demonstrate how state-space models can be adapted to financial applications where parameter stability cannot be assumed. Empirically, we document systematic patterns in how different investor classes respond to market conditions, with implications for understanding price discovery and market stability.
The remainder of this paper is organized as follows. Section 2 reviews related literature. Section 3 presents our methodological framework. Section 4 describes the data. Section 5 presents empirical results. Section 6 discusses robustness checks. Section 7 concludes.
@@ 2.1 Order Flow and Price Discovery
Empirical work has extensively documented this relationship across markets and frequencies (Hasbrouck,, 1991; Chordia et al.,, 2002; Lee and Ready,, 1991). Studies of investor heterogeneity find that different trader types exhibit distinct information content (Barber and Odean,, 2008; Griffin et al.,, 2003; Kaniel et al.,, 2008), with institutional investors generally exhibiting superior performance relative to retail traders.
@@ 2.2 Regime-Switching in Financial Markets
Regime-switching models have a long history in financial econometrics, originating with Hamilton, (1989) analysis of business cycle dynamics. Applications to asset returns include Ang and Bekaert, (2002) study of international equity correlations and Guidolin and Timmermann, (2006) analysis of portfolio allocation under regime uncertainty.
In market microstructure, Hasbrouck and Seppi, (2001) documents how liquidity conditions vary across market states, while Chordia et al., (2001) show that liquidity dries up precisely during stress periods when it is most needed. Our contribution extends this literature by explicitly modeling how order flow–return relationships shift across regimes.
@@ 2.3 Behavioral Asymmetries
Our asymmetric response framework connects to this literature by documenting differential reactions to positive and negative market shocks at the daily frequency, revealing that retail investors exhibit momentum-chasing behavior while foreign investors provide contrarian liquidity.
@@ 3.1 Adaptive Kalman Filter
We model the relationship between observed order flow and the “true” informed signal using a state-space representation. Let $S_{t}$ denote the observed market-cap normalized order flow at time $t$ and $\theta_{t}$ the latent informed signal. The state equation describes the evolution of the informed signal:
@@
Our key innovation is making the measurement noise variance $R_{t}$ heteroskedastic and explicitly coupled to market conditions:
@@ 3.2 Markov-Switching Regime Model
We specify a three-state Markov-switching model for market returns:
@@ 3.3 Asymmetric Response Function
To capture differential investor reactions to positive versus negative shocks, we estimate:
@@ 3.4 Integrated Strategy
We combine the three components into an “All-Weather” strategy:
@@ 4.2 Variable Construction
Following the literature on informed trading (Chordia et al.,, 2002), we construct market-cap normalized order flow as:
$S_{t}^{MC,(i)}=\frac{Buy_{t}^{(i)}-Sell_{t}^{(i)}}{MCAP_{t-1}}$ (8)
where $Buy_{t}^{(i)}$ and $Sell_{t}^{(i)}$ denote buy and sell volume in Korean Won for investor type $i$, and $MCAP_{t-1}$ is lagged market capitalization.
This normalization ensures that order flow is comparable across stocks of different sizes and has the interpretation of the fraction of market cap being net acquired by investor type $i$.
@@ 4.3 Summary Statistics
Table 1 presents summary statistics for our main variables. The average market-cap normalized foreign flow is slightly negative ($0.0023\%$), indicating net selling over the sample period. Individual investor flow is positive on average ($0.0084\%$), consistent with these investors being net buyers during the period. Daily returns average $0.03\%$ with substantial cross-sectional and time-series variation (standard deviation $3.5\%$).
@@ 5.1 Kalman Filter Performance
The Kalman filter provides consistent, albeit modest, improvements in predictive $t$-statistics and $R^{2}$ values. For foreign investors, the improvement in $t$-statistics ranges from $0.18\%$ at the 1-day horizon to $0.21\%$ at the 20-day horizon. The improvement increases with forecast horizon, suggesting the filter is more effective at extracting persistent information content.
@@
Figure 1 illustrates the relationship between Kalman gain and market volatility. The gain remains high (above 0.99) across all volatility conditions, indicating that the raw signals already contain substantial information content. However, the gain does decline slightly during high-volatility periods, consistent with the intended noise-reduction mechanism.
@@ 5.2 Regime Identification
Table 3 presents the characteristics of the three identified market regimes. The Bull regime (528 days, 43% of sample) features high daily returns ($0.154\%$), low volatility (0.54%), and exceptional risk-adjusted performance (Sharpe ratio 4.50). The Normal regime (598 days, 49%) exhibits near-zero returns ($0.034\%$) with moderate volatility (1.24%). The Crisis regime (95 days, 8%) shows strongly negative returns ($0.223\%$) and high volatility (3.87%).
@@
Bull 528 0.154% 0.54% 4.50 -0.0027%
Normal 598 -0.034% 1.24% -0.44 -0.0040%
Crisis 95 -0.223% 3.87% -0.92 0.0100%
The critical finding is the regime-conditional variation in foreign investor predictive power. During Bull periods, the foreign flow coefficient is $\beta=0.00023$ ($t=7.81$). This increases to $\beta=0.00064$ ($t=21.15$) in Normal regimes and rises dramatically to $\beta=0.00204$ ($t=19.23$) during Crisis periods—an 8.9-fold increase relative to Bull markets.
This pattern suggests that foreign investors possess (or reveal) superior information precisely when markets are most stressed. One interpretation is that during crises, noise trading by retail investors increases, making the relative information advantage of foreign institutions more pronounced.
@@
Figure 2 displays the time series of regime probabilities. The COVID-19 crash of March 2020 is clearly identified as a Crisis period, as are several subsequent stress episodes. The model identifies 130 crisis days (10.6% of the sample) with probability exceeding the 30% threshold.
@@ 5.3 Asymmetric Response Functions
Table 4 reports the asymmetric response coefficients by investor type. All three investor groups exhibit statistically significant asymmetric responses to market shocks (Wald test $p<0.001$ in all cases).
Table 4: Asymmetric Response Coefficients
Investor Type $\beta^{+}$ $t(\beta^{+})$ $\beta^{-}$ $t(\beta^{-})$ Ratio p-value
Foreign -0.000035*** (-12.02) 0.000070*** (22.60) -2.02 0.0000
Institutional -0.000021*** (-10.57) -0.000045*** (-20.24) 2.21 0.0000
Individual 0.000089*** (22.61) 0.000014*** (3.36) 0.16 0.0000
∗∗∗p<0.01, ∗∗p<0.05, p<0.10
Foreign investors display contrarian behavior: they sell after positive shocks ($\beta^{+}=-0.000035$) and buy after negative shocks ($\beta^{-}=+0.000070$). The asymmetry ratio of 2.02-2.02 indicates buying intensity after crashes is approximately twice the selling intensity after rallies.
Individual investors exhibit the opposite pattern: strong buying after positive shocks ($\beta^{+}=+0.000089$) and weak selling after negative shocks ($\beta^{-}=+0.000014$). The asymmetry ratio of $0.16$ reveals momentum-chasing behavior—retail investors “chase” rallies while being reluctant to sell during declines.

### v2

@@ Abstract
Most empirical microstructure research assumes that order flow–return parameters are constant, yet these relationships shift substantially across market regimes. Combining adaptive Kalman filtering, Markov-switching regime identification, and asymmetric response estimation, we characterize regime-dependent investor behavior in the Korean stock market during 2020–2024 using daily transaction data disaggregated by investor type. Three principal findings emerge: foreign investor predictive power increases several-fold during crisis periods relative to bull markets; individual investors chase momentum asymmetrically, reacting far more strongly to positive than to negative shocks; and independent information-theoretic validation corroborates both patterns. Rigorous out-of-sample testing reveals that these in-sample regularities do not generalize reliably, underscoring the need for proper validation methodology in microstructure research.
@@ 1 Introduction
While state-space filtering (Durbin and Koopman,, 2012), regime-switching models (Hamilton,, 1989), and asymmetric response estimation (Kahneman and Tversky,, 1979) each have long histories in financial econometrics, their joint application to investor-type order flow data has not been systematically explored for emerging markets during recent crisis and post-crisis periods. This paper addresses that gap for the Korean stock market during 2020–2024.
We make four contributions. First, we implement an Adaptive Kalman Filter that treats the “true informed signal” as a hidden state obscured by noise trading. We make the measurement noise variance heteroskedastic, explicitly coupling it to realized market volatility so that the filter automatically discounts high-volume panic trading as noise rather than signal. This provides a clean comparison of raw versus filtered order-flow predictors across volatility regimes.
Second, we embed investor-type order-flow/return relations in a Three-State Markov-Switching Model (Bull, Normal, Crisis), documenting how the predictive content of foreign, institutional, and retail flows varies across market regimes. This extends prior work on investor-type behavior and regime dependence in Korea (Choe et al.,, 2005, 1999; Kim and Wei,, 2002) by providing a unified characterization over the COVID and post-COVID period; Section 5.2 quantifies the regime-conditional variation.
Third, we estimate Asymmetric Response Functions that capture differential investor reactions to positive versus negative shocks. Consistent with the behavioral finance literature on Korean investors (Barber and Odean,, 2013), individual investors react to positive and negative shocks with markedly different intensity—a pattern we quantify in Section 5.3—while foreign investors provide contrarian liquidity during stress.
Fourth, we subject the integrated framework to rigorous out-of-sample validation with all parameters frozen at end-2022. The resulting failure to generalize (Section 5.5) underscores the importance of proper validation methodology in microstructure research.
The remainder of this paper is organized as follows. Section 2 reviews related literature. Section 3 presents our methodological framework, including the information-theoretic validation tools. Section 4 describes the data. Section 5 presents empirical results. Section 6 discusses robustness checks and cross-validates the framework using Jensen-Shannon Divergence and Viterbi decoding. Section 7 concludes.
@@ 2.1 Order Flow and Price Discovery
Empirical work has extensively documented this relationship across markets and frequencies (Hasbrouck,, 1991; Chordia et al.,, 2002; Lee and Ready,, 1991). Studies of investor heterogeneity find that different trader types exhibit distinct information content (Barber and Odean,, 2008; Griffin et al.,, 2003; Kaniel et al.,, 2008), with institutional investors generally exhibiting superior performance relative to retail traders. However, most of this work estimates unconditional coefficients, leaving open the question of how each trader type’s informational role varies across market regimes—a gap our analysis directly addresses.
@@ 2.2 Regime-Switching in Financial Markets
Regime-switching models have a long history in financial econometrics, originating with the analysis of business cycle dynamics by Hamilton, (1989). Applications to asset returns include the study of international equity correlations by Ang and Bekaert, (2002) and the analysis of portfolio allocation under regime uncertainty by Guidolin and Timmermann, (2006). Yet these applications have focused primarily on aggregate market dynamics rather than on how the informational content of specific investor types varies across regimes—a gap we address by embedding investor-type order flow within a regime-switching framework.
In market microstructure, Hasbrouck and Seppi, (2001) documents how liquidity conditions vary across market states, while Chordia et al., (2001) show that liquidity dries up precisely during stress periods when it is most needed. More recently, Kitamura, (2016) applies Markov-switching models to price impact dynamics and market crash identification, demonstrating that microstructure parameters exhibit substantial regime-dependence. We apply similar regime-dependent modeling to investor-type order flow in the Korean market, extending this literature to the COVID and post-COVID period.
@@ 2.3 Behavioral Asymmetries
Our asymmetric response framework connects to this literature by testing whether differential reactions to positive and negative market shocks at the daily frequency vary systematically across investor types in the Korean market.
2.4 Investor Behavior in Korean Markets
The Korean stock market provides uniquely detailed investor-type classification, enabling extensive research on behavioral differences across foreign, institutional, and individual investors. Choe et al., (2005) examine whether domestic investors have an informational edge relative to foreign investors in Korea, finding mixed evidence that varies by period and market conditions. Choe et al., (1999) analyze foreign investor behavior during the 1997 Korean crisis, documenting that foreigners engaged in positive-feedback trading that may have exacerbated market volatility during the crash. Kim and Wei, (2002) extend this analysis by comparing foreign portfolio investor behavior before and during crisis periods across multiple emerging markets, showing that crisis episodes fundamentally alter trading patterns and information revelation. These findings motivate our regime-switching framework, which tests whether the crisis-period shifts in foreign investor behavior documented in 1997 recur—and can be systematically quantified—during the COVID-19 episode.
More recently, Caporale et al., (2021) analyze how different investor types’ trading activity affects market volatility in Korea, finding that the relationship between investor flows and volatility is state-dependent and varies across investor classes. Broader studies of Korean retail investor behavior (Barber and Odean,, 2013) document systematic behavioral biases including overtrading, the disposition effect, and momentum-chasing in attention-grabbing stocks. The state-dependent volatility patterns identified by Caporale et al., (2021) provide direct motivation for our asymmetric response analysis, which decomposes these investor-volatility relationships into distinct positive- and negative-shock channels.
Our analysis updates this literature by characterizing investor-type order-flow/return relationships during the 2020–2024 period, encompassing both the COVID-19 crisis and the subsequent high-interest-rate environment, and by embedding these relationships within an integrated regime-switching and adaptive filtering framework.
@@ 3.1 Adaptive Kalman Filter
We model the relationship between observed order flow and the “true” informed signal using a state-space representation. Let $S_{t}$ denote the observed normalized order flow at time $t$ and $\theta_{t}$ the latent informed signal. The state equation describes the evolution of the informed signal:
@@
We adapt the standard Kalman filter by making the measurement noise variance $R_{t}$ heteroskedastic and explicitly coupled to market conditions (Durbin and Koopman,, 2012; Pole et al.,, 1994):
@@ 3.2 Markov-Switching Regime Model
Following the extensive literature on Markov-switching models in finance (Hamilton,, 1989; Ang and Bekaert,, 2002; Guidolin and Timmermann,, 2006; Kitamura,, 2016), we specify a three-state model for market returns:
@@ 3.3 Asymmetric Response Function
Building on the behavioral finance literature documenting asymmetric reactions to gains versus losses (Kahneman and Tversky,, 1979; Shefrin and Statman,, 1985; Odean,, 1998), we quantify differential investor reactions to positive versus negative shocks by estimating:
@@ 3.4 Integrated Strategy
We combine the three components into an “All-Weather” strategy that exploits the complementarity of the preceding methods:
@@
The rationale for this layered design is as follows. The Kalman filter (step 1) reduces measurement noise in raw order flow, providing a cleaner input to downstream decisions. Regime-conditional position sizing (step 2) addresses the finding that order flow coefficients vary by a factor of 3–5 across market states: rather than applying a single position size derived from unconditional estimates, the strategy scales exposure inversely with crisis-state probability, reducing risk when coefficient uncertainty is highest. Asymmetric stop-loss rules (step 3) incorporate the behavioral finding that positive and negative shocks trigger different investor reactions; stop-loss thresholds are set tighter for the direction in which momentum-chasing retail flow amplifies adverse moves. Together, these components translate the descriptive findings of Sections 3.1–3.3 into a testable portfolio framework evaluated in Section 5.4.
3.5 Information-Theoretic Validation Tools
We cross-validate the parametric framework above using two non-parametric information-theoretic tools.
Jensen-Shannon Divergence (JSD). For two discrete distributions $P$ and $Q$, the JSD is defined as:
$\text{JSD}(P\|Q)=\frac{1}{2}D_{KL}(P\|M)+\frac{1}{2}D_{KL}(Q\|M),\quad M=\frac{1}{2}(P+Q)$ (8)
where $D_{KL}$ is the Kullback-Leibler divergence. JSD is symmetric, bounded in $0,\ln 2]$ nats, and well-defined even when supports differ. We compute rolling JSD between each signal’s recent distribution and its long-run reference to construct adaptive blending weights that dynamically favor whichever normalization is more distributionally stable.
Viterbi Decoding. We estimate a two-state Gaussian Hidden Markov Model on four standardized order-flow features ($S^{TV}_{foreign}$, flow persistence, volatility ratio, turnover) at the daily market level. The Viterbi algorithm (Viterbi,, 1967) decodes the most likely state sequence, classifying each day as “Informed” (higher absolute signal, persistent flow) or “Noise” (lower signal, anti-persistent flow). This classification is independent of the three-state Markov-switching model in Section 3.2, enabling cross-validation.
@@ 4.2 Variable Construction
We construct normalized order flow signals using investor-type-specific normalizations motivated by the Matched Filter Framework (Kang,, 2025). This framework establishes that optimal signal extraction requires matching the normalization denominator to each investor type’s scaling behavior: investors whose order sizes scale with trading volume should use volume normalization, while those scaling with market capitalization should use market-cap normalization.
Table 14 validates these patterns in our sample through correlation analysis between absolute order flow $|D_{i}|$ and both market capitalization ($M$) and trading value ($V$). Foreign investors exhibit volume-scaling behavior ($\rho_{V}=0.66>\rho_{M}=0.59$) consistent with algorithmic execution (VWAP/TWAP). Domestic institutional investors show similar volume-scaling in raw correlations, but horse-race regressions (Table 15) reveal that market-cap normalization captures their true informational content ($t=15.26$) while volume normalization loses significance when both compete.
For foreign investors, we use trading value normalization:
$S_{t}^{TV,foreign}=\frac{Buy_{t}^{foreign}-Sell_{t}^{foreign}}{V_{t}}$ (9)
where $V_{t}$ is the daily trading value (volume $\times$ price). This normalization acts as a “matched filter” for volume-scaling traders, confirmed by horse-race regression ($t=50.85$).
For institutional investors, we use market-cap normalization:
$S_{t}^{MC,inst}=\frac{Buy_{t}^{inst}-Sell_{t}^{inst}}{MCAP_{t-1}}$ (10)
where $MCAP_{t-1}$ is lagged market capitalization. Horse-race regressions confirm this captures informed trading signals ($t=15.26$, positive coefficient).
For individual investors, we also use market-cap normalization despite their noise-trading nature. While individual investors’ underlying behavior is attention-driven, their order sizes scale with market capitalization ($\rho_{M}=0.61$) due to pronounced large-cap herding (Barber and Odean,, 2013). The $S^{MC}$ signal exhibits significant negative predictive power (t=15.01t=-15.01, $\beta=-0.28$), serving as a powerful contrarian indicator. This negative coefficient—documented in Section 5.3’s asymmetric response analysis—reflects momentum-chasing behavior that reliably predicts subsequent reversals.
@@ 4.3 Summary Statistics
Table 1 presents summary statistics for our main variables. The average TV-normalized foreign flow is slightly negative, indicating net selling over the sample period. Individual investor flow is positive on average, consistent with these investors being net buyers during the period. Daily returns average $0.03\%$ with substantial cross-sectional and time-series variation (standard deviation $3.5\%$).
@@ 5.1 Kalman Filter Performance
The Kalman filter provides consistent, albeit modest, improvements in predictive $t$-statistics and $R^{2}$ values (Figures 6 and 7 in Appendix A visualize the filtered versus raw signals and compare filter variants). For foreign investors, the improvement in $t$-statistics ranges from $0.18\%$ at the 1-day horizon to $0.21\%$ at the 20-day horizon. The improvement increases with forecast horizon, suggesting the filter is more effective at extracting persistent information content.
@@
To further validate our filtering approach, Table 3 compares the Kalman filter against alternative smoothing methods including moving averages (MA5, MA20, MA60) and exponentially weighted moving averages (EWMA10, EWMA30, EWMA50). The MA60 filter achieves the best Sharpe ratio (0.25) among all variants, though the raw signal produces the highest t-statistics (9.39). This suggests that longer-window smoothing may improve risk-adjusted returns despite reducing statistical significance.
Table 3: Comparison of Alternative Signal Filters
Filter t-stat R2 Sharpe Calmar Max DD
Raw 9.39∗∗∗ -0.1838 -0.55 -0.17 -53%
MA5 3.97∗∗∗ -0.1835 -0.43 -0.14 -53%
MA20 -0.30 -0.1822 -0.19 -0.08 -41%
MA60 -2.93∗∗∗ -0.1808 0.25 0.23 -18%
EWMA10 1.47 -0.1776 0.07 0.03 -41%
EWMA30 6.05∗∗∗ -0.1775 -0.40 -0.14 -50%
EWMA50 7.99∗∗∗ -0.1773 -0.39 -0.13 -52%
Notes: t-statistics from Fama-MacBeth regressions averaged across 1-, 5-, and 20-day horizons. Sharpe and Calmar ratios from daily long-short portfolios. *, **, *** denote significance at 10%, 5%, 1% levels.
Figure 1 illustrates the relationship between Kalman gain and market volatility. The gain remains high (above 0.99) across all volatility conditions, indicating that the raw signals already contain substantial information content. The gain does decline during high-volatility periods, consistent with the intended noise-reduction mechanism, but the overall improvement is modest, indicating that KRX daily order flow data is relatively clean at the aggregate level. The filter’s value lies less in dramatic signal improvement and more in providing a principled, volatility-adaptive framework that integrates naturally with the downstream regime-switching analysis.
@@ 5.2 Regime Identification
Table 4 presents the characteristics of the three identified market regimes. The Bull regime (528 days, 43% of sample) features high daily returns ($0.154\%$), low volatility (0.54%), and exceptional risk-adjusted performance (Sharpe ratio 4.50). The Normal regime (598 days, 49%) exhibits near-zero returns ($0.034\%$) with moderate volatility (1.24%). The Crisis regime (95 days by modal assignment, 8% of sample) shows strongly negative returns ($0.223\%$) and high volatility (3.87%).
@@
Bull 528 0.154% 0.54% 4.50 -0.8223%
Normal 598 -0.034% 1.24% -0.44 -0.8631%
Crisis 95 -0.223% 3.87% -0.92 1.0516%
Table 5 presents the regime-dependent price impact coefficients for all three investor types. The regime-conditional variation in foreign investor predictive power is striking. During Bull periods, the foreign flow coefficient is $\beta=0.00534$ ($t=25.04$). This increases to $\beta=0.00816$ ($t=38.32$) in Normal regimes and rises dramatically to $\beta=0.01873$ ($t=25.44$) during Crisis periods—a 3.51-fold increase relative to Bull markets.
Table 5: Regime-Dependent Price Impact Coefficients
Investor Type Bull Normal Crisis Ratio
[…]
(25.04) (38.32) (25.44)
[…]
(13.16) (13.22) (5.21)
[…]
(-9.64) (-17.54) (-14.85)
Notes: RAW (unstandardized) coefficients from daily cross-sectional regressions. Foreign investor flows normalized by trading volume; institutional and individual by market cap. t-statistics in parentheses. *, **, *** denote significance at 10%, 5%, 1% levels.
Institutional investors also show regime-dependent coefficients ($\beta=0.445$ in Bull, $0.619$ in Crisis, a 1.39-fold increase), but the variation is more moderate than for foreign investors. Individual investor flow carries negative predictive power across all regimes, with the magnitude rising sharply in Crisis periods ($\beta=-0.825$, t=14.85t=-14.85)—a 5.50-fold increase over Bull markets—consistent with intensified retail herding during stress.
This pattern extends prior evidence on foreign versus domestic investor behavior in Korean crisis episodes (Choe et al.,, 1999, 2005; Kim and Wei,, 2002) by quantifying the regime-specific magnitude of foreign informativeness during the COVID-19 period (see also Figures 8 and 11 in Appendix A). The finding suggests that foreign investors possess or reveal superior information precisely when markets are most stressed, plausibly because heightened retail noise trading during crises widens the gap between informed and uninformed flow.
@@
Figure 2 displays the time series of regime probabilities. The COVID-19 crash of March 2020 maps to a Crisis period, as do several subsequent stress episodes. Using a softer classification, 130 days (10.6% of the sample) carry Crisis-state probability exceeding 30%, reflecting partial regime overlap around transition dates.
@@ 5.3 Asymmetric Response Functions
Table 6 reports the asymmetric response coefficients by investor type (Figure 9 in Appendix A provides the full response curves). All three investor groups exhibit statistically significant asymmetric responses to market shocks (Wald test $p<0.001$ in all cases).
Table 6: Asymmetric Response to Market Shocks
Investor Type $\beta^{+}$ $\beta^{-}$ Ratio ($\beta^{-}/\beta^{+}$) NN Interpretation
Foreign 0.00172*** 0.00912*** 5.31 152,389 Contrarian
(10.72) (40.51)
Institutional -0.0000205*** -0.0000453*** 2.21 152,471 Symmetric
(-10.57) (-20.24)
Individual 0.0000888*** 0.0000142*** 0.16 152,471 Momentum
(22.61) (3.36)
[…]
Foreign investors display contrarian behavior: they respond more strongly to negative shocks ($\beta^{-}=0.00912$, $t=40.51$) than to positive shocks ($\beta^{+}=0.00172$, $t=10.72$). The asymmetry ratio of $5.31$ indicates that foreign flow response to market crashes is 5.3-fold stronger than their response to rallies, consistent with providing liquidity during stress periods.
Institutional investors fall between these extremes: both $\beta^{+}$ and $\beta^{-}$ are negative (indicating net selling after shocks of either sign), with a moderate asymmetry ratio of $2.21$. This near-symmetric rebalancing behavior is consistent with portfolio-driven trading rather than directional information or behavioral bias.
Individual investors exhibit the opposite pattern to foreign investors: strong response to positive shocks ($\beta^{+}=0.0000888$, $t=22.61$) and weak response to negative shocks ($\beta^{-}=0.0000142$, $t=3.36$). The asymmetry ratio of $0.16$ reveals momentum-chasing behavior—retail investors “chase” rallies at 6.3-fold the intensity of their response to declines. Figure 3 visualizes these contrasting response patterns across investor types.